Question 4 options:

FC.73 Ties at BYU-Idaho have become very popular since Tie the Knot, an IBC company, has started operations this semester. Sales have consistently increased during the weeks and the company wants to forecast future daily sales using the following data:

Day Week 1 Week 2 Week 3 Week 4 Week 5 Week 6 Week 7
Monday 32 29 37 32 35 28 36
Tuesday 22 22 26 30 21 31 30
Wednesday 17 19 21 23 20 25 24
Thursday 22 21 28 28 24 30 32
Friday 19 22 23 24 30 25 27

What is the seasonal index for Monday? (Display your answer to two decimal places.) 1.25

What is the seasonal index for Wednesday? (Display your answer to two decimal places.)   .81

What is the seasonal index for Friday? (Display your answer to two decimal places.) .93

Tie the Knot is projecting that sales for Week 8 will be 200 ties. Based on this forecast for Week 8, how many ties should they expect to sell on Tuesday in that week? (Display your answer to two decimal places.) 39.78

ANSWERS ARE IN BOLD ABOVE!
Answer #1

Answer:

1) What is the seasonal index for Monday?

- So, Seasonal Index for days = Total demand foa a day/ total demand

Day Demand for Monday is = 32+29+37+32+35+28+36 = 229

Total demand = Addintion of all day demand = 915

Seasonal index for Monday = 229/915 = 0.25

2) What is the seasonal index for Wednesday?

Day Demand for Wednesday is = 17+19+21+23+20+25+24

Total demand = Addintion of all day demand = 149

Seasonal index for Wednesday = = 149/915= 0.16.

3)

Day Demand for Friday is = 19+22+23+24+30+25+27=170

Total demand = Addintion of all day demand = 915

Seasonal index for Monday = 170/915 = 0.18

4) week 8 of the tuesday will be

daily demand of tueday= 22+22+26+30+21+31+30= 182

seasonal index for tueday = 182/915=0.20

week 8 of Tuesday = 200*0.20= 40

* Hope this is helpful to you, thank you & rate.